# Computing the Unit Group of a Commutative Finite $\mathbb{Z}$-Algebra

### v1

@@ Abstract.
1991 Mathematics Subject Classification:
Primary 13P99; Secondary 20-08, 11R27
@@ 3. Exponent Lattices in Finitely Generated $\mathbb{Z}$-Algebras
Let $R$ be a ring and $I$ an ideal in $R$. If $I:f^{\infty}=I:f^{m}$ some $f\in R$ and $m\in\mathbb{N}$. Then we have
@@
Let $R=P/I$ be a $\mathbb{Z}$-algebra, and let $f_{1},\dots,f_{k}\in R^{\times}$. The following sequence of instructions forms an algorithm which computes the exponent lattice of $(f_{1},\dots,f_{k})$ in $R$.
@@
It is an open question to what extent Algorithm 3 can be generalized to the non-commutative case. A straightforward generalization does not seem to be possible, since in the non-commutative case exponent lattices are in general not computable. This follows from the fact that the subgroup membership problem is undecidable for $4\times 4$ integral matrices (see [23]).
@@ 4. The Unit Group of Reduced Finite $\mathbb{Z}$-Algebras
This example demonstrates that we cannot directly reduce to the integral case. Instead we notice that if $R$ is torsion-free then $R$ is an order in the reduced zero-dimensional $\mathbb{Q}$-algebra $\mathbb{Q}\otimes_{\mathbb{Z}}R$.
@@
The unit group of an order in a zero-dimensional reduced $\mathbb{Q}$-algebra can be computed using the algorithm presented in Section 3 of [10]. In the following we present a modified version of this algorithm.
@@
Let $A$ be a reduced zero-dimensional $\mathbb{Q}$-algebra, and let $e_{1},e_{2}$ be orthogonal idempotents with $e_{1}+e_{2}=1$. Let $\mathcal{O}$ be an order in $A$. Consider the ideal $J=(e_{1}\mathcal{O}\cap\mathcal{O})+(e_{2}\mathcal{O}\cap\mathcal{O})$ and form the ring $S=\mathcal{O}/J$. Consider the ring homomorphisms $\varphi_{i}:e_{i}\mathcal{O}\rightarrow S$ given by $\varphi_{i}(e_{i}a)=a+J$ for $a\in\mathcal{O}$. Then we have
@@
Let $A$ be a reduced zero-dimensional $\mathbb{Q}$-algebra, and let $\mathcal{O}$ be an order in $A$. The following instructions form an algorithm which computes a set of generators of $\mathcal{O}^{\times}$.
@@
    The $\mathbb{Q}$-algebra $A=\mathbb{Q}\otimes_{\mathbb{Z}}R$ is zero-dimensional and its maximal ideals are given by $\mathfrak{m}_{i}=\mathfrak{p}_{i}\mathbb{Q}[x_{1},\dots,x_{n}]$ for $i=1,\dots,m$. We can therefore compute elements $q_{i}\in\bigcap_{i\neq j}\mathfrak{m}_{j}$ and $p_{i}\in\mathfrak{m}_{i}$ such that $q_{i}+p_{i}=1$. Then the residue classes $\bar{q}_{1},\dots,\bar{q}_{m}$ in $A$ are the primitive idempotents of $A$.
@@
To prove (c), we notice that the ideal $\bar{q}_{i}R\cap R$ is contained in the right-hand side, since $q_{i}\in\bigcap_{i\neq j}\mathfrak{m}_{j}$. To show the opposite inclusion, let $f\in P$ such that $\bar{f}\in\bigcap_{i\neq j}\mathfrak{p}_{j}/I$. Then we have $f=q_{i}f+p_{i}f$. Since $p_{i}\in\mathfrak{m}_{i}$, this implies $p_{i}f\in\bigcap_{i=1,\dots,m}\mathfrak{p}_{i}=I$. Hence we get $\bar{f}=\bar{q}_{i}\bar{f}\in\bar{q}_{i}R\cap R$. Part (d) follows analogously.
@@
Let $R=P/I$ be a reduced torsion-free finite $\mathbb{Z}$-algebra. The following instructions form an algorithm which computes a set of polynomials such that their residue classes generate $R^{\times}$.
@@
9:     Assuming $U=\{g_{1},\dots,g_{k}\}$, compute a set of generators $B\subseteq\mathbb{Z}^{k+\ell}$ of the exponent lattice of $\left(\bar{g}_{1},\dots,\bar{g}_{k},\bar{h}_{1}^{-1},\dots,\bar{h}_{\ell}^{-1}\right)$ in $P/J$.
@@
(Computing the Unit Group of a Reduced Finite $\mathbb{Z}$-Algebra) Let $R=P/I$ be a reduced finite $\mathbb{Z}$-algebra. The following instructions form an algorithm which computes generators of the unit group $R^{\times}$.
@@ 5. The Unit Group of Non-Reduced Finite $\mathbb{Z}$-Algebras
Using the results of the previous section, in particular Algorithm 4, we can now compute generators of the unit group of $R/\operatorname{Rad}(0)$. The next task is to lift these generators to a generating set of $R^{\times}$.
@@
Since we already saw how to determine generators of $R/\operatorname{Rad}(0)^{\times}$, it remains to compute generators of $1+\operatorname{Rad}(0)$. If $\operatorname{Rad}(0)^{2}=0$, then elements $1+f,1+g$ of $1+\operatorname{Rad}(0)$ satisfy $(1+f)(1+g)\equiv 1+f+g$. In this case generators of the additive group of $\operatorname{Rad}(0)$ immediately yield generators of $1+\operatorname{Rad}(0)$. If the nilpotency index of $\operatorname{Rad}(0)$ is greater than 2, we can inductively compute generators of $1+\operatorname{Rad}(0)^{i-1}$ in $R/\operatorname{Rad}(0)^{i}$.
@@
Let $R=P/I$ be a finite $\mathbb{Z}$-algebra. The following instructions form an algorithm which computes generators of the unit group $R^{\times}$.
@@
After line 3 has been executed, the list $G$ contains generators of the unit group of $R/\operatorname{Rad}(0)$. Now assume that the elements in $G$ generate $R/\operatorname{Rad}(0)^{i-1}$ after iteration number $i-1$ of lines 5–8. By Lemma 15, the group homomorphism
@@
The generators of $R/\operatorname{Rad}(0)^{i-1}$ together with the generators of $1+\operatorname{Rad}(0)^{i-1}$ therefore generate $\left(R/\operatorname{Rad}(0)^{i}\right)^{\times}$. Hence, after iteration number $s$, the list $G$ contains generators of $\left(R/\operatorname{Rad}(0)^{s}\right)^{\times}=R^{\times}$.
@@
Clearly, $\mathcal{H}$ is contained in $1+\operatorname{Rad}(0)$. By the proof of Algorithm 5, it is then enough to show that there are $g_{1},\dots,g_{s}\in H$ such that the4 residue classes $\bar{g}_{1},\dots,\bar{g}_{s}$ in $R/\operatorname{Rad}(0)^{i}$ generate $1+\operatorname{Rad}(0)^{i-1}$. Every element in $\operatorname{Rad}(0)^{i-2}$ can be written as a $\mathbb{Z}$-linear combination of the residue classes of the terms in $\mathcal{G}$. Thus, every element in $\operatorname{Rad}(0)^{i-1}$ can be written as a $\mathbb{Z}$-linear combination of the terms in $\left\{\bar{t}\bar{f}_{i}\mid i=1,\dots,k,\;t\in\mathcal{G}\right\}$. The claim now follows from the fact that $(1+\bar{f})(1+\bar{g})=1+\bar{f}+\bar{g}$ for $f,g\in\operatorname{Rad}(0)^{i-1}$, since $fg\in\operatorname{Rad}(0)^{i}$.
Let us apply Algorithm 5, together with this simplification, to a concrete example.
@@
Let $R=P/I$ be a finite $\mathbb{Z}$-algebra. The following instructions form an algorithm which computes the rank and the invariant factors of the group $R^{\times}$.
@@
4:  Let $r$ be the number of main diagonal entries of $S$ equal to zero, and let $k_{1},\dots,k_{u}$ be the non-zero main diagonal entries of $S$.

### v3

@@ Abstract.
2020 Mathematics Subject Classification. 68W30, 20F05, 13P99, 16Z05.
@@ 3. Exponent Lattices in Finitely Generated $\mathbb{Z}$-Algebras
Let $R$ be a ring, $I$ an ideal in $R$, and $f\in R$. If $m\in\mathbb{N}$ is such that $I:f^{\infty}=I:f^{m}$, then
@@
Let $R=P/I$ be a $\mathbb{Z}$-algebra, and let $f_{1},\dots,f_{k}\in R^{\times}$. Consider the following sequence of instructions.
@@
This is an algorithm which computes the exponent lattice of $(f_{1},\dots,f_{k})$ in $R$.
@@
It is an open question to what extent Algorithm 3 can be generalized to the non-commutative case. A straightforward generalization does not seem to be possible, since in this case the multiplicative relations between units are in general not computable. This follows from the fact that the subgroup membership problem is undecidable for $4\times 4$ integral matrices (see [23]). Also note that multiplicative relations in general do not form a lattice in the non-commutative case.
@@ 4. The Unit Group of Reduced Finite $\mathbb{Z}$-Algebras
This example demonstrates that we cannot directly reduce to the integral case. Instead we notice, that if $R$ is torsion-free, then $R$ is an order in the reduced zero-dimensional $\mathbb{Q}$-algebra $\mathbb{Q}\otimes_{\mathbb{Z}}R$.
@@
The unit group of an order in a zero-dimensional reduced $\mathbb{Q}$-algebra, can be computed using the algorithm presented in Section 3 of [10]. In the following we present a modified version of this algorithm.
@@
Let $A$ be a reduced zero-dimensional $\mathbb{Q}$-algebra and let $e_{1},e_{2}$ be orthogonal idempotents with $e_{1}+e_{2}=1$. Let $\mathcal{O}$ be an order in $A$. Consider the ideal $J=(e_{1}\mathcal{O}\cap\mathcal{O})+(e_{2}\mathcal{O}\cap\mathcal{O})$ and form the ring $S=\mathcal{O}/J$. Consider the ring homomorphisms $\varphi_{i}:e_{i}\mathcal{O}\rightarrow S$ given by $\varphi_{i}(e_{i}a)=a+J$ for $a\in\mathcal{O}$. Then we have
@@
Let $A$ be a reduced zero-dimensional $\mathbb{Q}$-algebra, and let $\mathcal{O}$ be an order in $A$. Consider the following sequence of instructions.
@@
This is an algorithm which computes a set of generators of $\mathcal{O}^{\times}$.
@@
    The $\mathbb{Q}$-algebra $A=\mathbb{Q}\otimes_{\mathbb{Z}}R$ is zero-dimensional and its maximal ideals are given by $\mathfrak{m}_{i}=\mathfrak{p}_{i}\mathbb{Q}[x_{1},\dots,x_{n}]$ for $i=1,\dots,m$. We can therefore compute elements $q_{i}\in\bigcap_{i\neq j}\mathfrak{m}_{j}$ and $p_{i}\in\mathfrak{m}_{i}$ such that $q_{i}+p_{i}=1$. The residue classes $\bar{q}_{1},\dots,\bar{q}_{m}$ in $A$ then form the primitive idempotents of $A$
@@
To proof (c), we notice that the ideal $\bar{q}_{i}R\cap R$ is contained in the right-hand side, since $q_{i}\in\bigcap_{i\neq j}\mathfrak{m}_{j}$. To show the opposite inclusion, let $f\in P$ such that $\bar{f}\in\bigcap_{i\neq j}\mathfrak{p}_{j}/I$. Then we have $f=q_{i}f+p_{i}f$. Since $p_{i}\in\mathfrak{m}_{i}$, this implies $p_{i}f\in\bigcap_{i=1,\dots,m}\mathfrak{p}_{i}=I$. Hence, $\bar{f}=\bar{q}_{i}\bar{f}\in\bar{q}_{i}R\cap R$. Part (d) follows analogously.
@@
Let $R=P/I$ be a reduced torsion-free finite $\mathbb{Z}$-algebra. Consider the following sequence of instructions.
@@
9:     Assume that $U=\{g_{1},\dots,g_{k}\}$, and compute a set of generators $B\subseteq\mathbb{Z}^{k+\ell}$ of the exponent lattice of $\left(\bar{g}_{1},\dots,\bar{g}_{k},\bar{h}_{1}^{-1},\dots,\bar{h}_{\ell}^{-1}\right)$ in $P/J$.
@@
This is an algorithm which computes a set of polynomials such that their residue classes generate $R^{\times}$.
@@
(Computing the Unit Group of a Reduced Finite $\mathbb{Z}$-Algebra) Let $R=P/I$ be a reduced finite $\mathbb{Z}$-algebra. Consider the following sequence of instructions.
@@
This is an algorithm which computes a generating set of the unit group $R^{\times}$.
@@ 5. The Unit Group of Non-Reduced Finite $\mathbb{Z}$-Algebras
Using the results of the previous subsection, in particular Algorithm 4, we can now compute generators of the unit group of $R/\operatorname{Rad}(0)$. The next task is to lift these generators to a generating set of $R^{\times}$.
@@
Since we already saw how to determine generators of $R/\operatorname{Rad}(0)^{\times}$, it remains to compute generators of $1+\operatorname{Rad}(0)$. If $\operatorname{Rad}(0)^{2}=0$, then elements $1+f,1+g$ of $1+\operatorname{Rad}(0)$ satisfy $(1+f)(1+g)\equiv 1+f+g$. In this case generators of the additive group of $\operatorname{Rad}(0)$, immediately yield generators of $1+\operatorname{Rad}(0)$. If the nilpotency index of $\operatorname{Rad}(0)$ is greater than 2, we can inductively compute generators of $1+\operatorname{Rad}(0)^{i-1}$ in $R/\operatorname{Rad}(0)^{i}$.
@@
Let $R=P/I$ be a finite $\mathbb{Z}$-algebra. Consider the following sequence of instructions.
@@
This is an algorithm which computes a generating set of the unit group $R^{\times}$.
@@
After line 3 has been executed the list $G$ contains generators of the unit group of $R/\operatorname{Rad}(0)$. Now assume that the elements in $G$ generate $R/\operatorname{Rad}(0)^{i-1}$ after iteration number $i-1$ of lines 5–8. By Lemma 15 the group homomorphism
@@
The generators of $R/\operatorname{Rad}(0)^{i-1}$ together with the generators of $1+\operatorname{Rad}(0)^{i-1}$ therefore generate $\left(R/\operatorname{Rad}(0)^{i}\right)^{\times}$. Hence, after iteration number $s$ the list $G$ contains generators of $\left(R/\operatorname{Rad}(0)^{s}\right)^{\times}=R^{\times}$.
@@
Clearly, $\mathcal{H}$ is contained in $1+\operatorname{Rad}(0)$. By the proof of Algorithm 5, it is then enough to show that there are $g_{1},\dots,g_{s}\in H$ such that $\bar{g}_{1},\dots,\bar{g}_{s}$ in $R/\operatorname{Rad}(0)^{i}$ generate $1+\operatorname{Rad}(0)^{i-1}$. Every element in $\operatorname{Rad}(0)^{i-2}$ can be written as a $\mathbb{Z}$-linear combination of the residue classes of the terms in $\mathcal{G}$. Thus, every element in $\operatorname{Rad}(0)^{i-1}$ can be written as a $\mathbb{Z}$-linear combination of the terms in $\left\{\bar{t}\bar{f}_{i}\mid i=1,\dots,k,\;t\in\mathcal{G}\right\}$. The claim now follows from the fact that $(1+\bar{f})(1+\bar{g})=1+\bar{f}+\bar{g}$ for $f,g\in\operatorname{Rad}(0)^{i-1}$, since $fg\in\operatorname{Rad}(0)^{i}$.
Let us apply Algorithm 5 using this simplification to a concrete example.
@@
Let $R=P/I$ be a finite $\mathbb{Z}$-algebra. Consider the following sequence of instructions.
@@
4:  Let $r$ be the number of diagonal entries of $S$ equal to zero, and let $k_{1},\dots,k_{u}$ be the non-zero diagonal entries of $S$.
@@
This is an algorithm which computes the rank and the invariant factors of the group $R^{\times}$.